Labco Scientific sells high-purity chemicals to universities, research laboratories and pharmaceutical companies. The company wants to invest in new equipment that will reduce shipping costs by better matching the size of the completed products with the size of the shipping container. The new equipment is estimated to cost $275,000 to purchase and install now. How much must Labco save each year for 5 years in order to justify the investment at an interest rate of 7% per year?
Answer #1

Solution -

We Know

Formula

A = P(A/P, i, n)


We need to find value of A,
P = $275,000; i = 7%; and n = 5 years
A = P(A/P, i, n)
(A/P, i, n) = 0.2439 ( from Factorial Table )
A = $275,000*0.2439

A= $67072.5

A= $ 67073 (Nearest Dollar)

So, Labco must save $ 67073 each year to purchase new equipment.

NOTE - DO UP VOTE . HAVE A NICE DAY